# Robust Bayesian Inference for Berkson and Classical Measurement Error Models

### v2

@@ 4.3 Special cases
Theorems 4.1 and 4.2 provide generalization error bounds in the general cases where no assumptions are made regarding the correct specification of the distributions of $X,X\>|\>W$ or $N$.
In many real-world cases (see e.g. Berry et al.,, 2002; Fung and Krewski,, 1999), the distribution of ME is known or can be estimated. In the Berkson error case, when $\operatorname{\mathbb{F}}^{0}_{N}$ is known or can be estimated, we can obtain a tighter bound for the generalization error, which vanishes as the data sample size grows large provided that we set $c$ large enough.
@@
Consider the special case when $\operatorname{\mathbb{F}}_{N}\equiv\operatorname{\mathbb{F}}_{N}^{0}$. Then, if $k^{2}_{X}$ is characteristic:
@@
$\quad\quad\quad\leq\frac{4}{\sqrt{n}}+2(1+\Lambda)\left(\sqrt{\frac{1}{c+2}}+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})\right).$ (25)
In the above Corollary, $k^{2}_{X}$ being characteristic is needed to ensure that MMD is a probability metric, i.e. $\operatorname{\operatorname{MMD}}_{k^{2}_{\mathcal{X}}}(\operatorname{\mathbb{P}},\operatorname{\mathbb{Q}})=0\iff\operatorname{\mathbb{P}}\equiv\operatorname{\mathbb{Q}}$ (Sriperumbudur et al.,, 2010). We note that if $\operatorname{\mathbb{F}}_{N}^{0}$ is known a priori, then it would make sense to set $c$ large so that the nonparametric posterior is centered around the well-specified prior. Taking the limiting case of $c\rightarrow\infty$, we observe that the upper bound of the generalization error approaches $\frac{2}{\sqrt{n}}$ and hence that the error decreases at a $\sqrt{n}$ rate. In the Classical error case, since the target joint distribution involves the marginal distribution of $X$ a similar assumption has to be made about the knowledge of $\operatorname{\mathbb{P}}_{X}^{0}$.
@@
Consider the special case when $\operatorname{\mathbb{F}}\equiv\operatorname{\mathbb{P}}_{X}^{0}$. Then, if $k^{2}_{X}$ is characteristic:
@@
$\quad\quad\quad\leq\frac{4}{\sqrt{n}}+2(1+\Lambda)\left(\frac{1}{\sqrt{n}(c+1)}+\sqrt{\frac{1}{c+2}}+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})\right).$ (26)
Another common assumption is that $N$ follows a Gaussian distribution (see f.ex. Cochran,, 1968; Spiegelman and Gray,, 1991). While not an assumption in our method, we explore this specific scenario when $k_{X}$ is the Gaussian kernel with parameter $l>0$. This allows for a closed form of the bound in Theorem 4.1 for $\mathcal{X}\subseteq\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}$.
@@
Let $k_{X}$ be the RBF kernel, $\operatorname{\mathbb{F}}_{N}^{0}=N(0,\sigma_{1}^{2}I)$ and $\operatorname{\mathbb{F}}_{N}=N(0,\sigma_{2}^{2}I)$. Then:
@@
$\quad\quad\leq\frac{4}{\sqrt{n}}+2(1+\Lambda)\left(\sqrt{\frac{1}{c+2}}+C_{1}\frac{c}{c+1}+\frac{C_{2}}{c+1}\right)$
where $C_{1}^{2}:=\left(\frac{l^{2}}{l^{2}+4\sigma_{1}^{2}}\right)^{d_{\mathcal{X}}/2}-2\left(\frac{l^{2}}{l^{2}+2\sigma_{1}^{2}+2\sigma_{2}^{2}}\right)^{d_{\mathcal{X}}/2}+\left(\frac{l^{2}}{l^{2}+4\sigma_{2}^{2}}\right)^{d_{\mathcal{X}}/2}$ and $C_{2}^{2}:=\left(\frac{l^{2}}{l^{2}+4\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}-2\left(\frac{l^{2}}{l^{2}+2\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}+1$.
@@
As expected, in the absence of ME, that is when $\sigma_{1}\approx 0$, we have $C_{2}\approx 0$. Similarly, we see that the closer $\sigma_{1}$ is to $\sigma_{2}$, the smaller $C_{1}$ is. This occurs because $\operatorname{\mathbb{F}}_{N}$ is closer to the true ME distribution $\operatorname{\mathbb{F}}_{N}^{0}$. Under the further assumption that the ME variance is known, i.e. $\sigma_{1}=\sigma_{2}$, as made by various existing methods (see Section 1), we have $C_{1}=0$.
5 Simulations & Applications
In this section, we investigate the empirical performance of our method for both Berkson and Classical ME. We first asses the robustness of our method to varying ME sizes in linear and nonlinear regression with synthetic data. Then, we examine two case studies in Econometrics and mental health science covering both Berkson and Classical error models.
5.1 Simulations
@@
We give an illustrative example of the performance of Robust-MEM with the TLS in Section 3.4.1. As is widely known (see e.g. Berkson,, 1950), Berkson ME does not generally lead to biased inference outcomes in linear regression so we will consider this simulation in the Classical error case. However, we note that there are cases where ignoring Berkson ME in linear regression would still introduce bias, for example in the presence of unobserved confounders as discussed in Haber et al., (2021). Consider the following linear regression model:
$y$ $=\theta_{1}+\theta_{2}x+\epsilon,\quad\epsilon\sim N(0,\sigma^{2}_{\epsilon})$
$w$ $=x+\nu,\quad\nu\sim N(0,\sigma^{2}_{\nu}).$
We sample $n=100$ points from a Gaussian distribution $x_{1:n}\sim N(10,4)$ and perform inference for the parameters $\theta_{0}=(\theta_{1},\theta_{2})=(1,2)$. We consider the Robust-MEM with TLS described in Section 3.4.1 and we further report results from frequentist Ordinary Least Squares (OLS) and TLS estimators as well as the SIMEX algorithm (Cook and Stefanski,, 1994) provided by the simex R package. In the Robust-MEM (TLS) algorithm We set $T=100$ and $c=1$ and vary $\sigma^{2}_{\nu}$ while keeping $\sigma_{\epsilon}$ fixed at 0.5 throughout this experiment. We want to examine the case where the ME distribution is not known exactly hence we use a prior centering measure $\operatorname{\mathbb{F}}_{N}=N(0,0.2)$ for all different true values of $\sigma_{\nu}^{2}$. We set the same value of apriori variance of $0.2$ in the SIMEX algorithm. Results for the well-specified case of error distribution are provided in Appendix C.1. The linear regression fit is illustrated in Figure 4 for all methods and the corresponding MSE and standard deviations of the parameter estimates over 100 runs are reported in Table 2. We observe that the Robust-MEM (TLS) method remains robust to ME even under a misspecified prior variance. In the absence of ME ($\sigma_{\nu}^{2}=1e-06$), OLS performs best whereas for small ME variance ($\sigma_{\nu}^{2}=0.5$) TLS achieves the smallest MSE with all methods performing well. The suggested method has the lowest MSE for higher values of the variance.
Figure 4: Model fit for linear regression in the misspecified prior case for an increasing ME variance (from left to right) over 100 simulations. The true model fit is shown along with our method (R-MEM (MMD)), nonlinear least squares (Least Squares) and SIMEX.
Method ME variance
1e-06 0.5 1.0 2.0
OLS $\theta_{1}$ 0.063 (0.083) 1.632 (1.232) 16.945 (7.153) 103.85 (21.629)
$\theta_{2}$ 0.001 (0.001) 0.016 (0.012) 0.17 (0.071) 1.035 (0.209)
R-MEM (TLS) $\theta_{1}$ 1.618 (0.128) 1.585 (0.161) 1.464 (0.222) 1.011 (0.335)
$\theta_{2}$ 0.013 (0.002) 0.013 (0.003) 0.012 (0.004) 0.011 (0.008)
SIMEX $\theta_{1}$ 0.07 (0.073) 1.127 (0.681) 15.404 (4.712) 99.524 (16.695)
$\theta_{2}$ 0.001 (0.001) 0.011 (0.007) 0.154 (0.046) 0.993 (0.159)
TLS $\theta_{1}$ 0.118 (0.156) 0.313 (0.561) 1.47 (1.948) 14.871 (11.412)
$\theta_{2}$ 0.001 (0.002) 0.003 (0.006) 0.015 (0.019) 0.15 (0.112)
Table 2: Mean Squared Error with associated standard deviation for $\theta=(\theta_{1},\theta_{2})$ in the Linear Regression example with Classical ME. Results are reported for an increasing ME variance ($\sigma^{2}_{\nu}$) in the observed data, over 100 simulations.
@@
We now consider an example where the mean regression function is nonlinear to illustrate the effectiveness of the MMD-based loss function discussed in Section 3.4.2. In the nonlinear case, directly addressing the TLS objective can result in inconsistent parameter estimates (Kukush and Zwanzig,, 2002; Van Huffel,, 2004). We consider a sigmoid-shaped curve defined by:
@@
with parameter of interest $\theta=(a,b)$. We generate $n=200$ observations with $\theta_{0}=(1,3)$ and $\sigma_{\epsilon}=0.5$. Further experimental details are provided in Appendix C. In the Classical error case, we compare again with SIMEX, which is readily available to use with Classical ME. In the Berkson case, we also report the frequentist minimum MMD estimator of Briol et al., (2019). This allows us to investigate how much of the robustness is gained by the NPL framework in comparison to the choice of loss function alone.
[…]
We start with the Berkson ME case in the misspecified case. The results in the well-specified case are similar and reported in Appendix C.2. The true ME distribution is Gaussian with increasing variance whereas the prior centering measure is set to be a Student-t distribution with 3 degrees of freedom. Results are reported for 100 simulations in Figure 5.
@@
Figure 5: Model fit for the sigmoid curve regression function with Berkson ME for an increasing ME variance (from left to right) over 100 simulations. The true model fit is shown along with our method (R-MEM (MMD)), nonlinear least squares (Least Squares) and minimum MMD estimator (MMD).
We observe that for small values of ME variance ($\leq 0.25$), all methods perform equally well, but as the variance increases the Robust-MEM (MMD) methodology remains the most robust to the ME. As previously mentioned, we wish to investigate the source of robustness from the Robust-MEM framework as compared to a frequentist minimum MMD estimator. To better study this, we report the associated MSE and their variances in table 3.
Method ME variance
$081e-$ 08 $0.25$ $1.0$ $4.0$
MMD $\theta_{1}$ 0.245 (0.503) 0.192 (0.503) 0.213 (0.219) 0.49 (0.335)
$\theta_{2}$ 1.335 (2.106) 0.953 (1.386) 1.526 (1.224) 3.846 (1.524)
R-MEM (MMD) $\theta_{1}$ 0.148 (0.257) 0.125 (0.226) 0.13 (0.179) 0.247 (0.293)
$\theta_{2}$ 0.19 (0.238) 0.271 (0.431) 0.431 (0.651) 0.769 (0.961)
Table 3: Mean Squared Error with associated standard deviation for $\theta=(\theta_{1},\theta_{2})$ in the Nonlinear Regression example with Berkson ME. Results are reported for an increasing ME variance ($\sigma^{2}_{\nu}$) in the observed data, over 100 simulations.
Compared to the frequentist minimum MMD estimator we can see that robustness is induced from the NPL framework and is not just a consequence of robustness properties of the MMD-based loss function. We further report the coverage of the obtained posteriors of our method in Table 4.
ME variance
1e-08 0.25 1 4
$\theta_{1}$ 0.86 0.92 0.94 0.87
$\theta_{2}$ 1 0.99 0.99 0.99
@@
Table 4: Coverage probability for the Robust-MEM (MMD) method in the Berkson error model with misspecified prior distribution for the ME. The coverage probability is calculated as the frequency over 100 simulations that the 90% posterior credible interval included the true parameter value.
[…]
We now evaluate the same model for the Classical ME case. We report results for both the well-specified and misspecified cases of prior centering measures. In the well-specified case in (11), we set $\operatorname{\mathbb{F}}_{X}:=N(0,1)$ and $\operatorname{\mathbb{F}}_{W\>|\>X}:=N(x,\sigma^{2}_{\nu})$ where $\sigma^{2}_{\nu}$ is the true ME variance. The prior centering measure $\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$ then has density $f_{X\>|\>w_{i}}(x\>|\>w_{i})\propto f_{W\>|\>X}(w_{i}\>|\>x)f_{X}(x)$ and is again a Normal distribution. In the misspecified case, we set $\operatorname{\mathbb{F}}_{X}:=N(0,1)$ and $\operatorname{\mathbb{F}}_{W\>|\>X}:=N(x,0.2)$. The obtained mean regression fit over 100 simulations is illustrated in Figure 6 along with the estimates obtained from standard nonlinear least squares and the SIMEX procedure.
(a) ME variance is correctly specified
(b) ME variance is misspecified
Method ME variance
$081e-$ 08 $0.25$ $1.0$ $4.0$
LS $\theta_{1}$ 0.096 (0.179) 0.077 (0.093) 0.118 (0.201) 0.139 (0.127)
$\theta_{2}$ 0.264 (0.551) 0.385 (0.314) 1.455 (0.491) 2.926 (0.36)
R-MEM (MMD) $\theta_{1}$ 0.1 (0.117) 0.101 (0.118) 0.106 (0.122) 0.113 (0.128)
$\theta_{2}$ 0.059 (0.078) 0.056 (0.079) 0.058 (0.087) 0.06 (0.09)
SIMEX $\theta_{1}$ 0.096 (0.179) 0.111 (0.202) 0.261 (1.562) 0.13 (0.127)
$\theta_{2}$ 0.264 (0.552) 0.323 (0.663) 0.748 (1.019) 2.244 (0.576)
Table 5: Mean Squared Error with associated standard deviation for $\bm{\theta}=(\theta_{1},\theta_{2})$ in the nonlinear regression example with Classical ME of known ME variance (well-specified case). Results are reported for an increasing ME variance ($\sigma^{2}_{\nu}$) in the observed data, over 100 simulations.
Method ME variance
$081e-$ 08 $0.25$ $1.0$ $4.0$
LS $\theta_{1}$ 0.096 (0.179) 0.077 (0.093) 0.118 (0.202) 0.139 (0.127)
$\theta_{2}$ 0.264 (0.551) 0.385 (0.314) 1.455 (0.491) 2.926 (0.36)
Robust-MEM (MMD) $\theta_{1}$ 0.1 (0.12) 0.137 (0.154) 0.252 (0.236) 0.642 (0.408)
$\theta_{2}$ 0.052 (0.063) 0.066 (0.08) 0.121 (0.134) 0.358 (0.295)
SIMEX $\theta_{1}$ 0.121 (0.27) 0.081 (0.093) 0.131 (0.31) 0.139 (0.128)
$\theta_{2}$ 0.46 (1.275) 0.337 (0.305) 1.414 (0.49) 2.919 (0.362)
Table 6: Mean Squared Error with associated standard deviation for $\bm{\theta}=(\theta_{1},\theta_{2})$ in the nonlinear regression example with Classical ME of unknown ME variance (misspecified case). Results are reported for an increasing ME variance ($\sigma^{2}_{\nu}$) in the observed data, over 100 simulations.
ME variance
1e-08 0.5 1.0 2.0
Well-specified $\theta_{1}$ 0.96 0.97 0.98 0.98
$\theta_{2}$ 1 1 1 1
Misspecified $\theta_{1}$ 0.96 0.97 0.98 0.97
$\theta_{2}$ 1 1 1 1
5.2 Berkson and Classical Measurement Error Studies
ME is especially important in health and social sciences applications in which contamination of the covariate observations can significantly bias the quantification of causal effects and consequently misguide decision-making. In this section, we explore two real-world datasets, one for each case of ME.

### v3

@@ 4.3 Special cases
Theorems 4.1 and 4.2 provide generalisation error bounds for any prior beliefs about the distributions of X,X|WX,X\>|\>W or $N$. In many real-world cases (see e.g. Berry et al.,, 2002; Fung and Krewski,, 1999), the distribution of ME is known or can be estimated. In the Berkson error case, when $\operatorname{\mathbb{F}}^{0}_{N}$ is known or can be estimated, we can obtain a tighter bound for the generalisation error, which vanishes as the data sample size grows large, provided that we set $c$ large enough.
@@
Let $\operatorname{\mathbb{F}}_{N}\equiv\operatorname{\mathbb{F}}_{N}^{0}$. Then, if $k_{X}$ is characteristic:
@@
$\quad\leq\frac{1}{\sqrt{n}}\left(2+\frac{2(1+\Lambda)}{\sqrt{c+2}}\right)+\frac{2c}{c+1}\sqrt{{\text{Var}}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{F}}_{N}^{0}]}$
$\quad+\frac{2}{c+1}\left(\Lambda\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})+\sqrt{2}\sqrt{\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})}\right).$
In the above corollary, $k_{X}$ being characteristic is needed to ensure that MMD is a probability metric, i.e. $\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{P}},\operatorname{\mathbb{Q}})=0\iff\operatorname{\mathbb{P}}\equiv\operatorname{\mathbb{Q}}$ (Sriperumbudur et al.,, 2010). We note that if $\operatorname{\mathbb{F}}_{N}^{0}$ is known a priori, then it would make sense to set $c$ large so that the nonparametric posterior is centered around the well-specified prior. Taking the limiting case of $c\rightarrow\infty$, we observe that the upper bound of the generalisation error approaches $\frac{1}{\sqrt{n}}\left(2+\frac{2(1+\Lambda)}{\sqrt{c+2}}\right)+2\sqrt{{\text{Var}}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{F}}_{N}^{0}]}$ and hence that the error decreases at a $\sqrt{n}$ rate with an additional error due to the RKHS variance of the true ME distribution. In the Classical error case, since the target joint distribution involves the marginal distribution of $X$ a similar assumption has to be made about $\operatorname{\mathbb{P}}_{X}^{0}$.
@@
Let $\operatorname{\mathbb{P}}_{X|w_{i}}\equiv\operatorname{\mathbb{P}}_{X}^{0}\forall i\in\{1,\dots,n\}$. Then, if $k_{X}$ is characteristic:
@@
$\quad\leq\frac{1}{\sqrt{n}}\left(\frac{2\Lambda}{c+1}+\frac{2\Lambda+2}{\sqrt{c+2}}\right)+\frac{2c}{c+1}\sqrt{{\text{Var}}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{P}}_{X}^{0}]}$
$\quad\quad+\frac{2}{c+1}\left(\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})+\sqrt{2}\sqrt{\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})}\right).$
Another common assumption is that the ME or the true covariate follows a Gaussian distribution (e.g. Spiegelman and Gray,, 1991). While not an assumption of our method, we explore this specific scenario when $k_{X}$ is the Gaussian kernel with parameter $l>0$. This allows for a closed form of the bound in Theorem 4.1 for $\mathcal{X}\subseteq\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}$.
@@
Let $k_{X}$ be the RBF kernel with lengthscale $l>0$, $\operatorname{\mathbb{F}}_{N}^{0}=\mathcal{N}(0,\sigma_{1}^{2}I)$ and $\operatorname{\mathbb{F}}_{N}=\mathcal{N}(0,\sigma_{2}^{2}I)$. Then,
@@
$\quad\quad\leq\frac{1}{\sqrt{n}}\left(2+\frac{2(1+\Lambda)}{\sqrt{c+2}}\right)+\frac{2c}{c+1}\left(\Lambda C_{1}+C_{2}\right)+\frac{2}{c+1}\left(\Lambda C_{3}+\sqrt{2}\sqrt{C_{4}}\right)$
where
$C_{1}^{2}$ $:=\left(\frac{l^{2}}{l^{2}+4\sigma_{1}^{2}}\right)^{\frac{d_{\mathcal{X}}}{2}}-2\left(\frac{l^{2}}{l^{2}+2\sigma_{1}^{2}+2\sigma_{2}^{2}}\right)^{\frac{d_{\mathcal{X}}}{2}}+\left(\frac{l^{2}}{l^{2}+4\sigma_{2}^{2}}\right)^{\frac{d_{\mathcal{X}}}{2}}$
$C_{3}^{2}$ $:=\left(\frac{l^{2}}{l^{2}+4\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}-2\left(\frac{l^{2}}{l^{2}+2\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}+1$
$C_{2}^{2}$ $:=1-2\left(\frac{l^{2}}{l^{2}+\sigma_{1}^{2}+\sigma_{2}^{2}}\right)^{\frac{d_{\mathcal{X}}}{2}}+\left(\frac{l^{2}}{l^{2}+2\sigma_{2}^{2}}\right)^{\frac{d_{\mathcal{X}}}{2}}$
$C_{4}^{2}$ $:=\left(\frac{l^{2}}{l^{2}+2\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}-2\left(\frac{l^{2}}{l^{2}+\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}+1.$
@@
As expected, in the absence of ME, that is when $\sigma_{1}\approx 0$, we have $C_{3}\approx 0$ and $C_{4}\approx 0$. Similarly, the closer $\sigma_{1}$ is to $\sigma_{2}$, the smaller $C_{1}$ and $C_{2}$ are. This occurs because $\operatorname{\mathbb{F}}_{N}$ is closer to the true ME distribution $\operatorname{\mathbb{F}}_{N}^{0}$. Under the further assumption that the ME variance is known, i.e. $\sigma_{1}=\sigma_{2}$, as made by various existing methods (see Section 1), we have $C_{1}=0$. In the Classical case, we get the following closed form when the true marginal distribution of $X$ is a Gaussian and the Gaussian priors are chosen for both the ME distribution and the marginal of $X$ as in the example in (9).
Corollary 4.2.4 (Classical).
Let $k_{X}$ be the RBF kernel with lengthscale $l>0$, $\operatorname{\mathbb{F}}_{N}^{0}=\mathcal{N}(0,\sigma_{1}^{2}I)$, $\operatorname{\mathbb{P}}_{X}^{0}:=\mathcal{N}(\mu_{X},\Sigma_{X})$ and $\operatorname{\mathbb{F}}_{X\mid w_{i}}:=\mathcal{N}(m_{i},\sigma_{\text{post}}^{2}I)$ obtained by the closed-form posterior in (9) using the prior $X\sim\mathcal{N}(m_{X},\sigma^{2}I)$ and $w_{i}\mid X\sim\mathcal{N}(X,\sigma^{2}_{N}I)$, i.e. $m_{i}:=\frac{\sigma^{2}}{\sigma^{2}+\sigma^{2}_{N}}w_{i}+\left(1-\frac{\sigma^{2}}{\sigma^{2}+\sigma^{2}_{N}}\right)m_{X}$ and $\sigma^{2}_{\text{post}}=\frac{\sigma^{2}\sigma^{2}_{N}}{\sigma^{2}+\sigma^{2}_{N}}$. Then,
$\quad\leq\frac{1}{\sqrt{n}}\left(\frac{2\Lambda}{c+1}+\frac{2\Lambda+2}{\sqrt{c+2}}\right)$
$\quad\quad+\frac{2c}{c+1}\left(\Lambda C_{1}+\frac{1}{n}\sum_{i=1}^{n}\sqrt{C_{2}+C_{i}^{2}}\right)+\frac{2}{c+1}\left(\Lambda C_{3}+\sqrt{2}\sqrt{C_{4}}\right)$
where
$C_{i}^{2}$ $:=\left|I+\tfrac{2\sigma_{\text{post}}^{2}}{\ell^{2}}I\right|^{-1/2}+\left|I+\tfrac{2}{\ell^{2}}\Sigma_{X}\right|^{-1/2}$
$\quad-2\left|I+\tfrac{1}{\ell^{2}}\big(\sigma_{\text{post}}^{2}I+\Sigma_{X}\big)\right|^{-1/2}\exp\!\left(-\tfrac{1}{2}\Delta_{i}^{\top}\big(\sigma_{\text{post}}^{2}I+\Sigma_{X}+\ell^{2}I\big)^{-1}\Delta_{i}\right)$
for $\Delta_{i}=m_{i}-\mu_{X}$,
$C_{2}$ $:=\left|I+\tfrac{2}{\ell^{2}}\Sigma_{X}\right|^{-1/2}$
$C_{1}^{2}$ $:=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\Big(1+\tfrac{4\sigma_{\text{post}}^{2}}{\ell^{2}}\Big)^{-d/2}\exp\!\left(-\tfrac{\|m_{i}-m_{j}\|^{2}}{4\sigma_{\text{post}}^{2}+\ell^{2}}\right)$
$\quad+\Big(1+\tfrac{4\sigma^{2}}{\ell^{2}}\Big)^{-d/2}-\frac{2}{n}\sum_{i=1}^{n}\Big(1+\tfrac{2(\sigma_{\text{post}}^{2}+\sigma^{2})}{\ell^{2}}\Big)^{-d/2}\exp\!\left(-\tfrac{\|m_{i}-m_{X}\|^{2}}{2(\sigma_{\text{post}}^{2}+\sigma^{2})+\ell^{2}}\right)$
and $C_{3},C_{4}$ are as in Corollary 4.2.3
Finally, under the assumption of RBF kernels, we prove that the generalisation error bounds presented in Theorems 4.1 and 4.2 are minimax optimal in terms of the rate in $n$ by employing Le Cam’s method (see Wainwright,, 2019, Lemma 15.9). Intuitively, this result follows from the fact that both bounds rely on a vanishing term of rate $\mathcal{O}(n^{-\frac{1}{2}})$ depending on the concentration rate of a minimum MMD estimator and two error terms which are independent of the sample size $n$ and dependent on the ME structure and the DP prior choice. As a result, we expect the $n$-rate to be at best the rate obtained in standard MMD estimation, without the added ME structure, which is $\mathcal{O}(n^{-1/2})$ (as in Briol et al.,, 2019; Chérief-Abdellatif and Alquier,, 2022; Dellaporta et al.,, 2022). Indeed, the minimax rate optimality of an MMD estimator under an RBF kernel has already been established by Tolstikhin et al., (2016). Here, we adapt this result to our setting and show minimax optimality of our generalisation error bounds in the case of an RBF kernel.
Proposition 4.2.1.
Let Assumption 4.0.2 hold and assume both $k_{X}$ and $k_{Y}$ are RBF kernels. Then for any inferential procedure that returns $\hat{\theta}_{n}$, given observations $\{(w_{i},y_{i})\}_{i=1}^{n}$ we have:
$\inf_{\hat{\theta}_{n}}\sup_{\operatorname{\mathbb{P}}\in\mathcal{P}_{\mathcal{X}\times\mathcal{Z}}}\operatorname{\mathbb{E}}_{(w_{i},x_{i},y_{i})_{1:n}\sim\operatorname{\mathbb{P}}}[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{\hat{\theta}_{n}})]\gtrsim n^{-\frac{1}{2}}.$
@@
5 Empirical Assesment
In this section, we empirically evaluate the performance of our method in the presence of Berkson and Classical measurement errors with synthetic data. Code to reproduce all experiments can be found at https://github.com/haritadell/me-models.
@@ 5 Empirical Assesment
We first examine the performance of Robust-MEM with the TLS loss. As is widely known (see e.g. Berkson,, 1950), Berkson ME does not generally lead to biased inference outcomes in linear regression so we consider the Classical error case. However, it’s worth noting that there are cases where ignoring Berkson ME in linear regression would still introduce bias, for example in the presence of unobserved confounders as discussed in Haber et al., (2021). Consider the linear regression model: $y=\theta_{1}+\theta_{2}x+\epsilon,\quad w=x+\nu$ where $\epsilon\sim N(0,\sigma^{2}_{\epsilon}),\quad\nu\sim N(0,\sigma^{2}_{\nu}).$ We sample $n=100$ points from a Gaussian distribution $x_{1:n}\sim N(10,4)$ and perform inference for the parameters $\theta_{0}=(\theta_{1},\theta_{2})=(1,2)$. In addition to Robust-MEM with the TLS loss, we include results from frequentist OLS and TLS estimators, as well as the SIMEX algorithm (Cook and Stefanski,, 1994) provided by the simex R package. We examine cases where the ME distribution is not precisely known, using a prior centering measure $\operatorname{\mathbb{F}}_{N}=N(0,0.2)$ for various true values of $\sigma_{\nu}^{2}$ and the same a priori variance of $0.2$ in the SIMEX algorithm. Results for a well-specified prior are provided in Appendix E.2. In Fig. 4 we observe that the Robust-MEM (TLS) method remains robust to ME even under a misspecified prior variance and high values of ME variance. In the absence of ME, OLS performs best with all methods performing well. More detailed results, including the MSE and variance can be found in Table 5 (Appendix E.2).
Figure 4: Fitted linear regression line over 100 simulations with a misspecified ME variance. The true line is shown along with (R-MEM (TLS)), nonlinear LS (Least Squares) and SIMEX.
@@
Directly addressing the TLS objective in nonlinear regression can result in inconsistent parameter estimates (Kukush and Zwanzig,, 2002; Van Huffel,, 2004) so we consider the Robust-MEM (MMD) with a sigmoid-shaped curve defined by:
@@
with parameter of interest $\theta=(a,b)$. We generate $n=200$ observations with $\theta_{0}=(1,3)$ and $\sigma_{\epsilon}=0.5$. We further consider the frequentist minimum MMD estimator of Briol et al., (2019) to investigate how much of the robustness is gained by the NPL framework in comparison to the choice of loss function alone. We consider Berkson ME where the true ME distribution is Gaussian with increasing variance whereas the prior centering measure is set to be a Student-t distribution with 3 degrees of freedom. The results for the well-specified prior and the Classical error cases are reported in Appendix E.3, E.4. We observe (Fig. 5 & Tab. 2) that for small values of ME variance ($\leq 0.25$), all methods perform satisfactorily, but as the variance increases Robust-MEM (MMD) remains the most robust to ME while also maintaining good coverage probability.
@@
Figure 5: Model fit for the sigmoid curve regression function with Berkson ME for an increasing ME variance over 100 simulations. The true model fit is shown along with our method (R-MEM (MMD)), nonlinear least squares (Least Squares) and minimum MMD estimator (MMD).
@@
Table 2: Nonlinear Regression example with Berkson ME (over 100 simulations) with a misspecified prior. Top: MSE with associated std. for $\theta=(\theta_{1},\theta_{2})$. Bottom: Coverage probability (based on the 90% posterior credible intervals) for the Robust-MEM (MMD) method.
Method/Metric ME variance
1e081e-08 $0.25$ $1.0$ $4.0$
MMD/ $\theta_{1}$ 0.25 (0.50) 0.19 (0.50) 0.21 (0.22) 0.49 (0.34)
MSE $\theta_{2}$ 1.34 (2.11) 0.95 (1.39) 1.53 (1.22) 3.85 (1.52)
R-MEM (MMD)/ $\theta_{1}$ 0.15 (0.26) 0.13 (0.23) 0.13 (0.18) 0.25 (0.29)
MSE $\theta_{2}$ 0.19 (0.24) 0.27 (0.43) 0.43 (0.65) 0.77 (0.96)
R-MEM (MMD)/ $\theta_{1}$ 0.86 0.92 0.94 0.87
Coverage prob. $\theta_{2}$ 1 0.99 0.99 0.99
This further demonstrates that robustness arises from the NPL framework rather than solely from the robustness properties of the MMD-based loss function.
Average Computational Times
In addition to the empirical performance results, we also recorded the average computational times for each method across our experiments. The experiments for the Robust-MEM and minimum MMD estimation methods were run on a GPU. The SIMEX method was run in R using the code provided by the simex R package. The average runtimes (in seconds) per dataset are summarised in Table 3.
Table 3: Average computational times (in seconds) with associated standard deviation across 100 simulation runs.
Method Linear Regression Nonlinear Regression
OLS 0.001 (0.0003)
TLS 0.003 (0.0008)
SIMEX 0.29 (0.03)
MMD 1.78 (0.23)
Robust-MEM (TLS) 5.41 (0.45)
Robust-MEM (MMD) 13.25 (3.52)
These results indicate that while the proposed Robust-MEM methods incur a higher computational cost relative to simpler approaches, such as OLS or TLS, the runtimes remain practical for the problem sizes considered due to the parallelisation property of the Posterior Bootstrap.
6 Berkson and Classical Measurement Error Studies
In this section, we examine the application of our framework in case studies from Econometrics and Mental Health Science in the presence of either Berkson or Classical ME.